OpenStudy (anonymous):

i have three problems i need figured out... PROBLEM 1: one draw is made from a deck of 52 cards. What is the probability that a king will be drawn? PROBLEM 2: A single ball is drawn from an urn containing two white, one red, and one blue ball. What is the probability that the blue ball will be drawn? i think the answer for this one is 1/4 but im not sure PROBLEM 3: a coin is tossed twice. What is the probability that it will show heads both times.?

OpenStudy (anonymous):

4 kings, 52 cards total 4/52 = 1/13

OpenStudy (anonymous):

#2 you have

OpenStudy (anonymous):

both head is \[\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\]

OpenStudy (anonymous):

thank you so much for explaining it to me it really helped a lot..... 8^)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

can you help me some more please